What is the equation of the line passing through (2, –3) and parallel to the line y = –6x – 1 in standard form?
Answer by lwsshak3(11628)

What is the equation of the line passing through (2, –3) and parallel to the line y = –6x – 1 in standard form?
***
Standard form of equation for a line: y=mx+b, m=slope, b=y-intercept
For given equation: y=-6x-1
slope,m=-6
slope of parallel line, m=-6
equation:
y=-6x+b
solve for b using coordinates of given point (2, -3)
-3=-6*2+b
b=9
Equation of parallel line: y=-6x+9